Jonas is 7 years older than Jeff. Last year Jonas was 12 years old. How old is Jeff now?
Answer by oscargut(2103)

Jonas is 7 years older than Jeff. Last year Jonas was 12 years old. How old is Jeff now?
Let x = Jonas`s age now
Let y = Jeff`s age now
x = y+7
x = 12+1 = 13
Then y = 6
Answer: 6
You can ask me more at: mthman@gmail.com
Thanks